Find the slope-intercept form of the equation of the line passing through the points:
(8, -6), (4, -2)

Answer by mananth(16946)

x1 y1 x2 y2
8 -6 4 -2

slope m = (y2-y1)/(x2-x1)
( -2 - -6 )/( 4 - 8 )
( 4 / -4 )
m= -1
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point (8,-6) in
Y = m x + b
-6.00 = -1 * 8 + b
-6.00 = -8 + b
b= -6 + 2
b= -8
So the equation will be
Y = -x+2
m.ananth@hotmail.ca